OpenStudy (gamereternity):

If f(x) varies directly with x, and f(x) = 6 when x = 20, which is the direct linear variation equation?

OpenStudy (gamereternity):

@John_ES

OpenStudy (gamereternity):

plz help

OpenStudy (gamereternity):

@Awolflover1 @AloneS @Agl202 @AnitaS. @ace-n-it @Aleah54

OpenStudy (gamereternity):

need answer

OpenStudy (gamereternity):

@bmk614 @Conqueror @chris215 @codyhelton75 @codyhelton75 @crazycocob

OpenStudy (gamereternity):

@kmcunningham

OpenStudy (gamereternity):

will fan and medal answers

rishavraj (rishavraj):

f(x) = kx when x = 20 f(x) = 6 get k plug it in the equation

OpenStudy (gamereternity):

?

OpenStudy (gamereternity):

fx=10/3x is that right

OpenStudy (gamereternity):

@John_ES is that right

OpenStudy (john_es):

It is just the inverse. 3/10

OpenStudy (gamereternity):

so f(x)=3/10x is the answer

OpenStudy (john_es):

\[f(x)=m\cdot x\Rightarrow 6=m\cdot 20\Rightarrow 6/20=m\Rightarrow 3/10=m.\]

OpenStudy (gamereternity):

ok ty

OpenStudy (john_es):

Yes, that is the answer.